# Measurements of decay branching fractions of the Higgs boson to hadronic final states at the CEPC

### v2

@@ 6 The determination of the branching fractions
Table 4: The measured branching fractions for the Higgs decays along with their statistical uncertainties are shown. The statistical uncertainty ranges from 0.58% ($H\rightarrow b\bar{b}$) to 9.81% ($H\rightarrow ZZ^{*}$).
Higgs boson decay $H\to b\overline{b}$ $H\to c\overline{c}$ $H\to gg$ $H\to WW^{*}$ $H\to ZZ^{*}$
branching fraction 0.5780 0.0292 0.0858 0.2150 0.0265
statistical uncertainty ±$\pm 0.58\%$$\pm 8.41\%$$\pm 2.99\%$$\pm 2.32\%$$\pm 9.81\%$
@@
$M^{T}_{mig}=\begin{pmatrix}\epsilon_{11}&...&\epsilon_{91}\\ ...&...&...\\ \epsilon_{19}&...&\epsilon_{99}\end{pmatrix}$ (3)
@@ 7 Results
In this analysis, by using the PFN method to separate events in $\mu^{+}\mu^{-}H$ process, the branching fractions of $H\to b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ at the CEPC, with a center-of-mass energy of 240 GeV and luminosity of 5600 fb-1, are measured to be 0.5780, 0.0292, 0.0858, 0.2150 and 0.0265, with the statistical uncertainty of $0.58\%$, $8.41\%$, $2.99\%$, $2.32\%$ and $9.81\%$, respectively.
@@
To account for the systematic uncertainty, the resolution of the detector was adjusted by increasing it by 2%. The resolution of particle transverse momenta is given as [7]:
$\sigma_{\frac{1}{p_{T}}}=2\times 10^{-5}\oplus\dfrac{1\times 10^{-3}}{p\sin^{3/2}\theta}\mathrm{GeV}^{-1}$ (5)
By applying the previous PFN model to MC samples generated with updated resolutions, the differences in branching fractions before and after the resolution change are considered as the systematic uncertainty. The systematic uncertainties for the branching fractions are estimated to be 1.29%, 9.69%, 0.67%, 2.61% and 1.11% for $b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ final states, respectively.
@@ 8 Conclusion
The Higgs boson branching fractions into $b\overline{b}/c\overline{c}/gg$ and $WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically, via the $Z(\mu^{+}\mu^{-})H$ process are studied using the PFN method at a center-of-mass energy of 240 GeV and a luminosity of 5600 fb-1 at the CEPC. Simulated samples of ”two-fermion” and ”four-fermion” processes are considered as backgrounds. The PFN model demonstrates strong performance in classifying different channels and generalizing across processes. The statistical uncertainty of branching fractions of $H\rightarrow b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ processes are estimated to be approximately $0.58\%$, $8.41\%$, $2.99\%$, $2.32\%$ and $9.81\%$, respectively. Compared to a previous analysis [17], which reported statistical uncertainties of $1.1\%$, $10.5\%$ and $5.4\%$ for the branching fractions of $H\to b\overline{b}/c\overline{c}/gg$ process, the PFN method achieves higher precision in a single execution, due to its better performance and deeper data exploitation. By increasing the transverse momentum resolution by 2% to account for differences between real data and simulated samples, the systematic uncertainties for the branching fractions are estimated to be 1.29%, 9.69%, 0.67%, 2.61% and 1.11% for $b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ final states, respectively. This study achieves highly precise measurements of decay branching fractions of Higgs, helping to increase the understanding of the properties of the Higgs boson and further testing of the Standard Model.
@@
The authors express sincere gratitude to Prof. Gang Li, Prof. Manqi Ruan and Prof. Yanping Huang for invaluable guidance and support throughout the research. The authors present special thanks to Dr. Fábio L. Alves, Dr. Jinfei Wu and Dr. Mingxu He for useful discussion and advice. The authors thanks the IHEP Computing Center for its firm support.

### v3

@@ 6 The determination of the branching fractions
Table 4: The measured branching fractions for the Higgs decays along with their statistical uncertainties are shown. The statistical uncertainty ranges from 0.55% ($H\rightarrow b\bar{b}$) to 15.81% ($H\rightarrow ZZ^{*}$).
Higgs boson decay $H\to b\overline{b}$ $H\to c\overline{c}$ $H\to gg$ $H\to\tau\overline{\tau}$ $H\to WW^{*}$ $H\to ZZ^{*}$
branching fraction 0.5770 0.0291 0.0857 0.0632 0.2150 0.0264
statistical uncertainty ±$\pm 0.55\%$$\pm 8.59\%$$\pm 3.03\%$$\pm 2.85\%$$\pm 1.58\%$$\pm 15.81\%$
@@
$M^{T}_{mig}=\begin{pmatrix}\epsilon_{1,1}&...&\epsilon_{12,1}\\ ...&...&...\\ \epsilon_{1,12}&...&\epsilon_{12,12}\end{pmatrix}$ (7)
@@ 7 Results
In this analysis, by using the PFN method to separate events in $\mu^{+}\mu^{-}H$ process, the branching fractions of $H\to b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ at the CEPC, with a center-of-mass energy of 240 GeV and luminosity of 5600 fb-1, are measured to be 0.5770, 0.0291, 0.0857, 0.0632, 0.2150 and 0.0264, with the statistical uncertainty of $0.55\%$, $8.59\%$, $3.03\%$, $2.85\%$, $1.58\%$ and $15.81\%$, respectively.
@@
To account for the systematic uncertainty, the resolution of transverse momentum of the detector was adjusted by increasing it by 2% to represent for differences between real data and simulated samples. By applying the previous PFN model to MC samples generated with updated resolutions, the differences in branching fractions before and after the resolution change are considered as the systematic uncertainty. The systematic uncertainties for the branching fractions are estimated to be 0.21%, 3.88%, 2.74%, 1.39%, 0.18% and 19.09% for $b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ final states, respectively.
@@ 8 Conclusion
The Higgs boson branching fractions into $b\overline{b}/c\overline{c}/gg$ and $\tau\overline{\tau}/WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically, via the $Z(\mu^{+}\mu^{-})H$ process are studied using the PFN method at a center-of-mass energy of 240 GeV and a luminosity of 5600 fb-1 at the CEPC. Simulated samples of ”two-fermion” and ”four-fermion” processes are considered as backgrounds. The PFN model demonstrates strong performance in classifying different channels and generalizing across processes. The statistical uncertainty of branching fractions of $H\rightarrow b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ processes are estimated to be approximately $0.55\%$, $8.59\%$, $3.03\%$, $2.85\%$, $1.58\%$ and $15.81\%$, respectively. Compared to a previous analysis [17], which reported statistical uncertainties of $1.1\%$, $10.5\%$ and $5.4\%$ for the branching fractions of $H\to b\overline{b}/c\overline{c}/gg$ process, the PFN method achieves higher precision in a single execution, due to its better performance and deeper data exploitation. By increasing the transverse momentum resolution by 2% to account for differences between real data and simulated samples, the systematic uncertainties for the branching fractions are estimated to be 0.21%, 3.88%, 2.74%, 1.39%, 0.18% and 19.09% for $b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ final states, respectively. This study achieves highly precise measurements of decay branching fractions of Higgs, helping to increase the understanding of the properties of the Higgs boson and further testing of the Standard Model.
@@
We would like to The authors present special thanks to CEPC higgs physics working group for useful discussion and advice. The authors thank the IHEP Computing Center for its firm support. This work is supported in part by National Key R&D Program of China under Contracts No. 2022YFE0116900 and the Basic Science Center Program No. 12188102 by National Natural Science Foundation of China.